Accelerator Update for March 9, 2009 - March 11, 2009 |
Monday March 9 The day shift began with Operations monitoring the Tevatron (TeV) store 6880 (130.23E30) and the Recycler stash (75.3E10), with the Antiproton Source stacking (26.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. At 10:03 AM, a Linac quadrupole power supply (L:QPS114) tripped off. It reset. At 11:02 AM, a Pbar power supply expert reported that A:IB was making a lot of noise. Operators transferred beam out of the Accumulator and then let the expert begin his inspection. The Pbar power supply expert completed his work on A:IB at 11:47 AM. At 1:18 PM, Operations tried to transfer antiprotons to the Recycler, but due to two files not matching, no transfer was made. A controls expert investigated. Operations reported at 1:53 PM that the MI Ion Profile Monitor (IPM) was repaired. A controls expert reported at 1:56 PM that she had rebooted the Pbar and Recycler shot transfer engines. This allowed Operations to complete a stash. Operations reported at 2:31 PM that over the next 30 minutes: Booster RF experts would bring BRF18 online, Linac experts would check the quadrupole timing, cryo system experts would tighten a Pbar dry engine belt, Booster experts would check the chopper “on” tube, and CUB experts would bleed the Booster 95 LCW pump. Operations resumed sending beam to Pbar, MiniBooNE, and NuMI at 3:08 PM. Operations reported at 4:01 PM that the Recycler had stashed antiprotons six times during the day shift. At 5:26 PM, a Linac Klystron RF station (KRF1) tripped off. It reset. Operations terminated store 6880 at 7:31 PM Operations put the TeV into shot setup at 7:43 PM. Operations established store 6881 at 9:01 PM with an initial luminosity of 286.11E30. At 11:29 PM, the Pbar sweeping magnet power supply (M:USWA) tripped off. It reset. Operations reported at midnight that the Recycler had stashed seven times during the evening shift. Tuesday March 10 The midnight shift began with Operations monitoring store 6881 (117.87E30) and the stash (147.1E10), with Pbar stacking (17.4E10), and with NuMI and MiniBooNE taking beam. At 12:56 AM, LRF3 tripped off. It reset. At 4:44 AM, the 750 KeV area buncher tripped off. It reset. At 5:17 AM, a Linac quadrupole power supply (L:QPS102) tripped off for the first of two times in about 75 minutes. It reset. At 6:13 AM, Operations reported having problems with a Linac quadrupole power supply (L:Q43). It would occasionally turn off the Linac beam permit. Operators investigated, but could find nothing wrong. At 6:15 AM, the 400 MeV area beam permit tripped off for the first of three times in about two hours due to bad beam positions. It reset. At 7:40 AM, L:Q43 again caused problems. At 7:51 AM, a CUB expert reported that an outside contracting crew, working in CUB, hit a Linac 55 LCW line and caused a leak. Operations contacted Linac experts. Operations reported at 7:54 AM that the Recycler had stashed nine times during the midnight shift. At 8:06 AM, the 400 MeV area tripped off for the first of three times in five minutes. It reset. CUB experts reported at 8:16 AM that they had switched to pump #5 for the Linac 55 LCW. A Pbar expert began a cooling rate study at 8:46 AM. Operations terminated store 6881 at 9:36 AM. TeV experts began a study at 9:47 AM. At 10:27 AM, A:IB was making noise again. Operations transferred antiprotons out of the Accumulator so a power supply expert could investigate. Operations prepared Pbar for access at 10:56 AM. The Pbar power supply expert reported at 11:44 AM that he had completed his work on A:IB. Pbar resumed stacking at 12:14 PM. At 3:19 PM, a Linac bulk power supply (L:Q6BULK) tripped off. Experts investigated. Operators put Pbar into standby. TeV experts completed their study at 3:56 PM. Operations prepared D0 and CDF for access at 3:57 PM. Operations reported at 4:06 PM that the Recycler had stashed once during the day shift. Operators reported at 4:15 PM that they had pumped out water from transformer pits at TeV sectors B2, C2, and at MI-52. Cryo system experts reported at 4:25 PM that they had finished their TeV sector A1 work and were beginning maintenance on the sector B3 system. At 4:45 PM, a Linac Bulk power supply (L:Q7BULK) tripped off. It reset. At 4:52 PM, LRF3 tripped off. It reset. CDF came out of access at 5:39 PM. D0 came out of access at 5:48 PM. At 6:47 PM, Operations reported that the TeV sector A1 cold compressor was showing a high motor temperature alarm. Cryo system experts investigated. At 6:53 PM, L:QPS114 tripped off. Operators swapped control cards, but that did not help. They put the original cards back and reset the supply. Cryo system experts reported at 7:22 PM that they had completed their TeV sector B3 wet engine work. Cryo system experts reported at 8:14 PM that they had repaired the TeV sector A1 cold compressor. Operations put the TeV through a dry squeeze at 9:12 PM. Operations put the TeV through a wet squeeze at 9:55 PM. Operations conducted a TeV orbit check at 10:03 PM. Operations put the TeV through another wet squeeze at 10:38 PM. Operations put the TeV into shot setup at 11:02 PM. At 11:14 PM, Operations reported that the Pbar abort kicker #1 alarmed. It reset. At 11:19 PM, L:QPS117’s current had dropped by 3 amps and then later by 5 amps. It eventually returned to its normal setting. Wednesday March 11 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (6.9E10), with Pbar stacking (168.9E10), and with NuMI and MiniBooNE taking beam. Operations established store 6888 at 12:26 AM with an initial luminosity of 302.14E30. At 5:05 AM, the NuMI beam permit tripped off when a power supply (E:V100) drifted low. Everything reset. At 6:06 AM, Operations reported a number of MI aborts due to bad beam positions. Operators lowered the beam intensity. Operations reported at 7:57 AM that the Recycler had stashed six times during the midnight shift. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow a TeV end-of-store study. |
|
More Information Comments and Suggestions |
| last modified 03/11/2009 email Fermilab |
| Security, Privacy, Legal | |